From: NASA HQ/SpaceRef/NASA Watch
Posted: Thursday, June 8, 2006
SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2006) to enhance access to related status reports and NASA activities.
All ISS systems continue to function nominally, except those noted previously or below.
CDR Pavel Vinogradov serviced the BMP (Harmful Impurities Removal System), starting the "bake-out" cycle to vacuum on absorbent bed #2 of the regenerable dual-channel filtration system. Before sleep time today, the bake-out will be terminated. [Regeneration of each of the two cartridges takes about 12 hours and is conducted only during crew awake periods. After the recent successful IVA/EVA installation of a new H2 dump line for the Elektron, the BMP now uses its vacuum vent line/valve without having to share it with the electrolysis machine.]
Elektron is off, having shut down offnominally last night after running for several hours on 24 amps without apparent issues. Cause of the shutdown is currently unknown (suspect: the power supply), and troubleshooting by RSC-Energia will get underway tonight with diagnostic system tests. [There is no immediate impact, since oxygen represses can still be performed from either Progress 20P or 21P O2 tanks.]
FE/SO Jeff Williams had two hours to do the periodic thorough inspection of the Emergency Lighting Power Supply (ELPS) units in the U.S. segment (USOS). [There are three ELPS units in the Node, two in the Lab, and one ELPS in the US Airlock.]
Both crewmembers in turn took their second periodic on-orbit hearing assessment (O-OHA) test, a NASA environmental health systems examination to assess the efficacy of acoustic countermeasures, using a special MEC (Medical Equipment Computer) application. [The O-OHA audiography test involves minimum audibility measurements for each ear over a wide range of frequencies (0.25-10 kHz) and sound pressure levels, with the crewmembers using individual-specific Prophonics earphones, Bose ANC headsets and the SLM (sound level meter). To conduct the testing, the experimenter is supported by special EarQ software on the MEC (Medical Equipment Computer), featuring an up/down-arrow-operated slider for each test frequency that the crewmember moves to the lowest sound pressure level at which the tone can still be heard. The baseline test is required not later than about Flight Day 14 for each new Expedition and is then generally performed once per month.]
The FE ran his daily atmospheric status check for ppO2 (Partial Pressure Oxygen) and ppCO2 (pp Carbon Dioxide), using the CSA-O2 (Compound Specific Analyzer-Oxygen Sensor) and the CDMK (CO2 Monitoring Kit).
Jeff also deployed two passive FMK (formaldehyde monitoring kit) sampling assemblies in the Lab (below CEVIS) and SM (most forward handrail), to catch any atmospheric formaldehyde on a collector substrate for subsequent analysis on the ground.
Vinogradov and Williams worked through the regular fire drill/OBT (on-board training), a mandatory periodic one-hour exercise specifically written for the current two-person crew. Primary goal of this exercise is to provide the station residents with the most realistic emergency training possible. The drill is always conducted with the support of both MCCs in close coordination. [OBT objectives are to (a) practice fire response procedures (FRPs) and all incorporated actions for the case of a software-detected fire to locate, extinguish, and verify extinguishing attempts; (b) browse through RS laptop and the Signal-VM fire detection system displays as well as the automated software (algorithms) response to the fire event; (c) practice crew communication necessary to perform emergency FRPs; (d) update the locations of support hardware (CSA-CP compound specific analyzer-combustion products, IPK-1M gas masks and OSP-4 fire extinguishers to be used for fire suppression in the FGB. These exercises do not actually use any fire equipment but simulate such actions to the maximum extent possible. The OBT concluded with a 15-min. debrief with Russian/U.S. ground specialists.]
The CDR had another 2.5 hrs. for cleaning up after last week's spacewalk by stowing EVA tools and updating the IMS (Inventory Management System) databases to reflect the transfers and locations.
As part of periodic preventive maintenance, Vinogradov replaced two OKR-1 portable fire extinguishers with new ones in the SM and one in the DC-1 Docking Compartment. [The replaced units were stowed for disposal on Progress.]
Pavel brought the new ESA experiment CULT up to date by filling out its "cultural" questionnaire on the RSE1 laptop. Ground specialists were available to tag up via S-band as necessary. [CULT is a study conducted currently by Russia for ESA. The multi-Increment investigation, which eventually will involve 12 subjects, is dedicated to the study of cultural aspects and leadership styles of on-board crews as a function of mission duration, including interactions within multinational crews. The questionnaire is contained on a PCMCIA memory card, to be used for all subjects and sessions.]
Jeffrey began another round of research with the BCAT-3 (Binary Colloidal Alloy Test 3) science experiment, now focusing on Sample 1. [Activities led off today with historical documentation and video setup, checking proper configuration and operation of the EarthKAM hardware at the MWA (Maintenance Work Area), Sample 1 initialization/homogenizing and run. The EK camera is taking automated time-lapse flash photography (once every hour) of BCAT sample #1 at the MWA (Maintenance Work Area), controlled from the SSC-7 laptop (Station Support Computer 7). The current run extends over four days, with Jeff conducting daily checks of the alignment and focus of the camera and position of flash aimed at the sample of the BCAT-3 science activity.]
The FE completed the regular weekly audit/inventory of the available CWCs (collapsible water containers) and their contents, to keep track of onboard water supplies. [Updated "cue cards" based on Jeff's water calldowns are sent up every other week. The current card lists 16 water containers (~180 liters total) for the four types of water identified on board: technical water (for Elektron, flushing, hygiene; one CWC was found leaking, #1040), potable water (~180 liters), condensate water (for processing) and other (ITCS contingency fluid, ITCS/EMU waste water). Current assumed rate of water use is 2.2 liters per person per day with Elektron (vs. 1.7 liters per person per day without Elektron). Updated resupply rate for Expedition 12, as of 3/27, was 1.53 liters per person per day. Total water currently on board is ~1197 liters, which would last ~272 days without resupply. This includes 420 l delivered by Progress 21. 22P is expected to bring 130 l and ULF1.1 ~645 l.]
The FE performed the regular monthly maintenance of the TVIS treadmill, mainly consisting of an inspection of the Russian and U.S. tie-down harnesses (straps & buckles) and associated SBS (Series Bungee System) for any damage. TVIS can currently not be used for exercise, but the crew has alternatives for the necessary aerobic exercise. [During the inspection, a tool (3/32" L-wrench) used to check belt tension, slipped out of Jeff's hand and into the TVIS chassis. In order to allow retrieval of the wrench, the treadmill belt will have to be de-tensioned. Until that time, the TVIS is off limits for the crew.]
Williams also did the periodic (once per month) routine inspection of the RED (Resistive Exercise Device) with canister cords, squat harness components, and accessory straps, and the canister bolts for re-tightening if required.
Both crewmembers worked out in their regular 2.5-hr. physical exercise program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE), RED (FE) and VELO bike with bungee cord load trainer (CDR). [Pavel Vinogradov's daily protocol prescribes a strict four-day microcycle exercise with 1.5 hr on the treadmill in unmotorized mode and one hour on VELO plus load trainer (today's TVIS run was not conducted, see above).]
Later today, Jeff will transfer the exercise data files to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).
Jeffrey also performed the daily routine maintenance of the Service Module (SM)'s environment control & life support system (SOZh), including the toilet system (ASU), and the weekly inspection of the air/liquid condensate separator apparatus (BRPK), and he also updated/edited the standard IMS "delta file", including locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur). [The BRPK is part of the condensate water processor (SRVK) that processes the condensate coming from the currently active air conditioner (SKV) for the Elektron.]
A major new task added yesterday to Pavel's discretionary "time permitting" work list is the onboard search for ~25 "missing" Russian equipment items, listed on an uplinked new 13-page log. [The list includes such items as a transformer for the SOZh's water management system with cable and accessories that were delivered on Progress 11 but never used, an air duct section with acoustic screen, a Rodnik pump unit, pressurization equipment, etc.]
As of 6/4, the crew had a total of 282.6 food rations on board (71.1 U.S., 211.5 Russian). Usage rate is one ration per crewmember per day. STS-121/ULF1.1 will deliver an additional 246 rations (82 container-equivalents). A list was uplinked to the crew last night which provided a preview of fresh food items to be delivered on ULF1.1 (and quite probably made their mouths water). [Included in the crew-requested groceries are such goodies as fresh fruit (apples, grapefruit, Kiwi, oranges, small onions, etc.), peanut butter, ketchup and mayo, BBQ sauce packets, pickets, salsa, double strength coffees (Vienna & French Vanilla Nut), etc. There will also be "bonus" items for the newly arriving third ISS crewmember, Thomas Reiter.]
The 4B1 battery capacity test begun in 6/6 was completed yesterday and is currently being evaluated. [NiH (nickel hydrogen) batteries can develop and display memory loss resulting in a temporary loss of capacity that can periodically be erased by fully discharging and charging cells (reconditioning). The battery state of charge (SOC) reported in telemetry does not include the effect of this reduced capacity. Reconditioning was performed on all P6 Battery sets starting in Oct. 2004 through Dec. 2005. In the current round, reconditioning of the first set, 2B2, was successfully finished on 4/12, followed by 4B3 & 4B2 on 5/23-24. These tests are necessary to improve battery health and to determine the amount of amp hour capacity retained since the reconditioning was performed.]
The four-day CMG-1 (Control Moment Gyroscope #1) wheel speed characterization test by MCC-H specialists ends today. [Purpose of the test, without crew involvement, was to finish the checkout of the new CMG-1 by commanding 16 different wheel speeds over the course of four days and monitoring the performance of the gyro package.]
The station continues in LVLH XVV (local vertical local horizontal/x-axis in velocity vector) attitude, as required by the current high solar Beta angle magnitude (31.0 deg), until the reboost and the maneuver to LVLH YVV (y-axis in velocity vector) tomorrow afternoon.
MCC-H specialists have developed a new automated control program for conditioning the BGA (Beta Gimbal Assembly) bearings, expected to be executed beginning on 6/12 while the station is in the YVV attitude. [Over the years, the BGAs have exhibited anomalous behavior off and on, resulting in motor stall trips. A variety of techniques have been implemented to mitigate this problem. The leading cause is thought to be a build-up of friction in the bearings over time which leads to a stall event. However, the natural motion that the arrays follow while tracking the sun in the sun-oriented XPOP attitude (x-axis perpendicular to orbit plane) has been shown to alleviate this friction and mitigate the problem. The new automated conditioning program is intended to imitate the alleviative features of that motion even while in bearing-stressing YVV.]
Tomorrow's reboost (2:47pm EDT), by 21P's DPO rendezvous & docking thrusters, followed by maneuver to YVV at 2:54pm, will be preceded by attitude control handover to RS at 12:30pm and concluded by its return to USOS at 3:45pm. [Designed to set up proper orbit phasing for 22P (docking 6/26) and ULF1.1/STS-121 (docking 7/3 earliest, with phasing remaining valid for the duration of the July launch window), the maneuver will apply a delta-V of 0.7 m/s to boost mean altitude by 1.3 km, with a burn duration of 2 min 58 sec.]
Today's CEO (crew earth observations) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, were Tunis, Tunisia (looking a touch left for Tunis which has a major airport. Although not part of the CEO city-margin mapping effort, Tripoli, Libya's capital city, lied just right of track ~2 minutes later. Major changes are taking place in and around both cities as Libya's oil reserves begin to be tapped at a greater pace), and Roter Kamm (red comb) Impact Crater, Namibia (looking right for this small, young and well-formed impact crater (2.5 km diam., 3.7 million years old). Visual cue: a line of dark sandstone outcrops on the track side of the target).
To date, more than 198,000 of CEO images have been taken in the first five years of the ISS, almost one third of the total number of images taken from orbit by astronauts.
CEO photography can be viewed and studied at the websites:
See also the website "Space Station Challenge" at:
To view the latest photos taken by the expedition 13 crew visit:
Expedition 13 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/
Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.
ISS Orbit (as of this morning, 8:15am EDT [= epoch]):
Significant Events Ahead (all dates subject to change):
ISS Altitude History
Apogee height -- Mean Altitude -- Perigee height
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://spaceflight.nasa.gov/realdata/sightings/ on NASA's Human Spaceflight website. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.
// end //